# Personalizing Agent Privacy Decisions via Logical Entailment

### v2

@@ 5.2.2. Privacy Preferences-based Baseline
We provide privacy judgments on prior requests made by a user. There are two prompts that we then consider.
@@
    ICL (w Undet). We modify the instruction in ICL to output ’undetermined’ if the model is unable to confidently judge whether the incoming request is appropriate or inappropriate based on the user’s privacy preferences. These account for scenarios where the user’s privacy preferences may not always provide sufficient information to accurately judge the incoming request. The prompts are in A.4.
@@
5.3. Models and Evaluation metrics
Models
We evaluate ARIEL and the baselines on three advanced models: Gemini 2.5-Pro (gemini-2.5-pro), GPT-5 (gpt-5-2025-08-07), and Claude Sonnet 4.5 (claude-sonnet-4-5-20250929). Additionally, we include a smaller, open source model, Gemma 3 4B (team2025gemma).
Metrics
We compare the agent’s predicted judgments against the sampled users’ judgments on incoming requests. We report separate F1 scores for both the appropriate and inappropriate classes. Given an evaluation dataset $D$ containing incoming requests $R_{u}$ with judgments $l_{u}$ for each user $u$, we define the precision ($P_{A}$) for the appropriate class ($c^{+}=\texttt{appropriate}$) as:
We also define recall ($R_{A})$ for the appropriate class as:
Then we can define the F1 score ($F1_{A}$) for the appropriate class as:
$F1_{A}=2\cdot\frac{P_{A}\cdot R_{A}}{P_{A}+R_{A}}.$
A similar process can be done to obtain the F1 Score for the inappropriate class $F1_{I}$.
Evaluation Procedure for general norms. For each unique request in the datasets, we generate a prompt as described in Section 5.2.1. This results in 825 prompts for the SPA dataset and 1,411 for the Education dataset, covering both zero-shot and privacy norms. Once the prompts are generated, we provide all of them to the models to obtain their responses. Then, for each user, we compare their judgment on an incoming request with the LLM’s judgment on that same request, evaluating this on both general norms prompt types.
Evaluation Procedure for user privacy preferences. We generate a prompt for each incoming request with the user’s privacy judgments on prior requests. This results in 5,000 prompts for the SPA dataset and 3020 prompts for the Education dataset, covering ICL, ICL (w Undet), and ARIEL. Incoming requests resulting in an ’undetermined’ judgment are excluded from the evaluation for the respective baselines (i.e., ICL (w Undet) or ARIEL). To account for this partial coverage, we report the Support (Sup) metric, defined as the total number of appropriateness judgments (i.e., ’appropriate’ or ’inappropriate’) made by the model.
6. Results
We present experiments, ablation studies, and error analysis to study the efficacy and robustness of ARIEL. First we evaluate the end-to-end performance improvement of ARIEL over existing baselines in Section 6.1. Then we evaluate how varying the number of prior user judgments affect ARIEL in Section 6.2. Finally, we qualitatively analyze the failure modes of ARIEL in Section 6.3.
@@ 6.1. End-to-End Evaluation
(1) ARIEL enhances personalized privacy judgments compared to pure LLM reasoning. ARIEL consistently outperforms LLM-based baselines across all datasets and reasoning models, with notable gains in appropriate judgments. For example, ARIEL improves upon ICL (w Undet) by 10.4% $F1_{\text{A}}$ with GPT-5 on the Education dataset. In other words, the $F1_{\text{A}}$ error reduces by 39.1%, demonstrating that logical entailment effectively grounds judgments in prior user privacy judgments. Notably, these improvements are magnified in smaller models. With Gemma 3 4B, ARIEL yields $F1_{\text{A}}$ gains of 13% and 17% over ICL (w Undet) on the SPA and Education datasets, respectively. This demonstrates ARIEL is suitable for on-device deployment of personalized agents.
Figure 6. Ablation study on the number of prior requests with user judgments contained in the user’s knowledge base $D_{u}$. We evaluate ICL w (Undet) and ARIEL on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class, and the total number of appropriateness judgments (Sup). We find that ARIEL is more robust to varying the number of prior requests compared to ICL w (Undet)
Figure 7. Ablation study on CoT. We evaluate zero-shot, privacy norms, and ICL w (Undet) on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class. The results suggest that generally CoT has negligible impact on privacy judgments across both models.
(2) ARIEL accurately identifies when the appropriateness of an incoming request is challenging to judge. We observe that the number judgments that were undetermined by ARIEL is higher than ICL (w Undet). For Gemini 2.5 Pro, we see that ARIEL judged 1203 and 1339 incoming requests as undetermined for SPA and Education datasets, respectively, compared to just 387 and 695 for ICL (w Undet). This observation raises the question of whether the appropriateness label of the incoming requests left undetermined by ARIEL can actually be inferred by an LLM based on the user’s prior privacy judgments.
@@
(3) Relying only on the LLM’s awareness of privacy norms is insufficient for personalization. We observe that performance on appropriate judgments from zero-shot is low for both models and datasets. For inappropriate judgments, zero-shot prompting does performs much better than on appropriate judgments, but the results remain suboptimal. Hence, the LLM’s awareness of privacy norms cannot sufficiently capture the nuanced behavior of user privacy preferences within the SPA and Education contexts.
@@
(4) General privacy norms extracted from overall user responses for each dataset can improve performance over zero-shot. We observe that privacy norms improve appropriate judgments for the SPA dataset across all models, and inappropriate judgments for the Education dataset on most. However, including general norms can sometimes degrade performance. For instance, with Gemini 2.5 Pro on the Education dataset, $F1_{\text{A}}$ decreases by $10\%$ while $F1_{\text{I}}$ improves by only $8\%$, suggesting that adding general privacy norms is not always beneficial.
@@
(5) Prior user privacy judgments improve personalized privacy judgments over general privacy norms. By comparing the $F1_{\text{A}}$ and $F1_{\text{I}}$ between ICL and both zero-shot and privacy norms for all datasets and models, it is clear that personalizing LLMs with prior user privacy judgments improves appropriate and inappropriate judgments over general privacy norms. The improvement over general privacy norms can in some cases be as large as 20% for $F1_{\text{A}}$.
(6) Personalized privacy judgments improve when including undetermined as an option. When comparing ICL and ICL (w Undet), we observe an improvement in $F1_{\text{A}}$ and $F1_{\text{I}}$ for all state-of-art reasoning models. These results suggest that a hybrid approach of combining automated agent judgments with occasional escalation to the user can be highly effective for personalization.
6.2. Ablation Evaluation
We evaluate how the number of prior request affect the performance of ARIEL and baseline approaches. Moreover, we investigate how the choice of reasoning affects LLM privacy judgments.
6.2.1. Number of prior requests
We explore how varying the number of prior requests with user judgments affect the agent’s judgments. Specifically, we compare the $F1_{A}$, $F1_{I}$, and Support (Sup) of ARIEL against ICL (w Undet) on the SPA dataset using Gemini 2.5 Pro. As evident in Figure 6, while the $F1_{I}$ gap remains relatively constant, the disparity in $F1_{A}$ widens significantly as the number of prior requests decreases. Notably, when reducing the prior requests from 60 to 20, $F1_{A}$ drops by only 3.4% for ARIEL, compared to an 8.8% decline for ICL (w Undet). This demonstrates that ARIEL is more robust to limited prior privacy judgments. Nonetheless, we observe that the number of entailments generated by ARIEL decreases substantially as the number of prior requests decreases. This result is expected, as there are fewer candidate requests available to satisfy the entailment criteria (i.e., fewer prior requests differ by exactly one parameter from the incoming request).
6.2.2. Assessing impact of reasoning with Chain of Thought (CoT)
To assess the impact of reasoning strategy choice on appropriateness judgments, we implement CoT by prompting the LLM to provide a reasoning before making a judgment (wei2022chain). Inspired by findings that CoT affects privacy leakage (mireshghallah2024can), we evaluate its effect on judgment accuracy across zero-shot, privacy norms, and ICL (w Undet). Figure 7 indicates that CoT typically provides negligible changes in performance. However, we observe for zero-shot with Gemini 2.5 Pro that the $F1_{A}(\%)$ and $F1_{I}(\%)$ is significantly impacted by CoT, both positively and negatively.
While CoT provides mixed results for general privacy norms, the results clearly demonstrate that adding CoT to both ICL and ICL (w Undet) consistently hurts their performance. When manually inspecting the generated reasoning traces, we notice that the LLM makes misaligned assertions about the user’s privacy judgments, which propagate all the way to the model’s final judgment (see Section 6.3 for more information). This result also demonstrates that eliciting explainability from LLM-based reasoning can degrade the appropriateness judgment. In contrast, ARIEL provides a clear, traceable reasoning while maintaining well-grounded personalized privacy judgments.
6.3. Qualitative Analysis
6.3.1. Error Analysis on incorrect appropriateness judgments for ICL (w Undet) + CoT
Analysis of the reasoning traces reveals that Gemini 2.5 Pro occasionally performs implicit entailment when identifying prior requests that share identical parameters except for one. This behavior is noteworthy, as we did not explicitly instruct the model to use this logic, but only to extract patterns from user privacy judgments. We find that the majority of errors stem from preference misalignment (14 examples), where the model draws incorrect equivalencies between distinct parameter values to do this entailment. For instance, given prior requests involving ’close friends’ or ’parents’ as recipients and an incoming request involving ’housemates,’ the model infers that the privacy judgment should remain consistent due to perceived similarity in type of recipient. However, the user may not view these three recipients as similar.
We also attribute a subset of errors to the model’s inherent conservatism regarding information sharing, which we describe as default to inappropriate (5 examples). This occurs when the model concludes it cannot use the most similar prior requests that are judged as ’appropriate’ by the user. But rather than outputting ’undetermined’, the model reasons that the judgment is inappropriate. Finally, we identified one instance of contradictory prior requests, where the provided user history contained judgments that logically conflicted with the ground truth for the incoming request.
6.3.2. Error analysis on incorrect appropriateness judgments for ARIEL
@@
We classify errors in ARIEL into three categories: (1) ontology generation, (2) ontology mapping, and (3) contradictory prior requests. The majority of errors stem from ontology generation, where the generated ontology produces incorrect entailments (10 examples). This typically arises from data sparsity when specific parameter values (e.g., ’housemates’ or ’close friends’) appear only once or twice in the user’s prior requests. The model struggles to accurately reflect the user’s privacy preferences onto the ontology. A second source of error is ontology mapping, where the model fails to map parameter values to the ontology correctly (5 examples). This is often caused by domain-specific idiosyncrasies in the evaluation dataset. For instance, the transmission principle has specific format combining both purpose and condition, which can be challenging for the model to map. Finally, the remaining errors (5 examples) result from contradictory prior requests. For instance, there was a prior request where user had judged sharing bank account details with business and finance skills for the purpose of making transactions as inappropriate. The incoming request is identical as the prior, except there is no purpose. Thus, the incoming request should also be inappropriate, since removing the purpose likely makes the request more inappropriate. Both ICL (w Undet) +CoT and ARIEL judged this request inappropriate, however, the user judged the incoming request to be appropriate.
@@ 7. Discussion
Based on the results and takeaways, we discuss in more details the broad implications of this work as well as its limitations.

### v3

@@ 5.2.2. Privacy Preferences-based Baseline
We provide privacy judgments on prior requests made by a user. We consider two prompts.
@@
    ICL (w Undet). We modify the instruction in ICL to output ’undetermined’ if the model is unable to confidently judge whether the incoming request is appropriate or inappropriate based on the user’s privacy preferences. This modification accounts for scenarios where the user’s privacy preferences may not always provide sufficient information to accurately judge the incoming request. The prompts are in A.4.
5.3. Models and Evaluation metrics
Models
We evaluate ARIEL and the baselines on three advanced models: Gemini 2.5-Pro (gemini-2.5-pro), GPT-5 (gpt-5-2025-08-07), and Claude Sonnet 4.5 (claude-sonnet-4-5-20250929). Additionally, we include a smaller, open-source model, Gemma 3 4B IT (team2025gemma).
Metrics
We compare the agent’s predicted judgments against each sampled user’s judgments on incoming requests. We report separate F1 scores for both the appropriate and inappropriate classes. Given an evaluation dataset $D$ containing incoming requests $R_{u}$ with judgments $l_{u}$ for each user $u$, we define the precision $P_{A}$ for the appropriate class ($c^{+}=\texttt{appropriate}$) as:
We also define recall $R_{A}$ for the appropriate class as:
Then we can define the F1 score $F1_{A}$ for the appropriate class as:
$F1_{A}=2\cdot\frac{P_{A}\cdot R_{A}}{P_{A}+R_{A}}.$
A similar process can be used to obtain the F1 score for the inappropriate class $F1_{I}$. High $F1_{A}$ indicates the agent is correctly sharing user data (utility), whereas high $F1_{I}$ indicates the agent is correctly withholding data (privacy).
Evaluation Procedure for general norms. For each unique request in the datasets, we generate a prompt as described in Section 5.2.1. This results in 825 prompts for the SPA dataset and 1411 for the Education dataset, covering both zero-shot and privacy norms. Once the prompts are generated, we provide all of them to the models to obtain their responses. Then, for each user, we compare their judgment on an incoming request with the LLM’s judgment on that same request across both general norms prompt types.
Evaluation Procedure for user privacy preferences. We generate a prompt for each incoming request with the user’s privacy judgments on prior requests. This results in 5000 prompts for the SPA dataset and 3020 prompts for the Education dataset, covering ICL, ICL (w Undet), and ARIEL. Incoming requests resulting in an ’undetermined’ judgment are excluded from the evaluation for the respective baselines (i.e., ICL (w Undet) or ARIEL). To account for this partial coverage, we report the Support (Sup) metric, defined as the total number of appropriateness judgments (i.e., ”appropriate” or ”inappropriate”) made by the model.
6. Results
@@
We present experiments, ablation studies, and error analysis to study the efficacy and robustness of ARIEL. First, we evaluate the end-to-end performance improvement of ARIEL over existing baselines in Section 6.1. Then, we evaluate how varying the number of prior user judgments affects ARIEL in Section 6.2. Finally, we qualitatively analyze the failure modes of ARIEL in Section 6.3.
@@ 6.1. End-to-End Evaluation
(1) ARIEL enhances personalized privacy judgments compared to pure LLM reasoning. ARIEL consistently outperforms LLM-based baselines across all datasets and reasoning models, with notable gains in appropriate judgments. For example, ARIEL improves upon ICL (w Undet) by 10.4% $F1_{\text{A}}$ with GPT-5 on the Education dataset. In other words, the $F1_{\text{A}}$ error reduces by 40.6%, demonstrating that logical entailment effectively grounds judgments in prior user privacy judgments. Notably, these improvements are magnified in smaller models. With Gemma 3 4B IT, ARIEL yields $F1_{\text{A}}$ gains of 13% and 17% over ICL (w Undet) on the SPA and Education datasets, respectively. This demonstrates ARIEL is suitable for on-device deployment of personalized agents.
(2) ARIEL accurately identifies when the appropriateness of an incoming request is challenging to judge. We observe that the number of judgments that were undetermined by ARIEL is higher than ICL (w Undet). For Gemini 2.5 Pro, we observe that ARIEL judged 1203 and 1339 incoming requests as undetermined for SPA and Education datasets, respectively, compared to just 387 and 695 for ICL (w Undet). This observation raises the question of whether the appropriateness label of the incoming requests left undetermined by ARIEL can actually be inferred by an LLM based on the user’s prior privacy judgments.
@@
(3) Relying only on the LLM’s awareness of privacy norms is insufficient for personalization. We observe that performance on appropriate judgments from zero-shot is low for both models and datasets. For inappropriate judgments, zero-shot prompting can perform much better than on appropriate judgments, but the results remain suboptimal. Hence, the LLM’s awareness of privacy norms cannot sufficiently capture the nuanced behavior of user privacy preferences within the SPA and Education contexts.
@@
(4) General privacy norms extracted from overall user responses for each dataset can improve personalized privacy judgments. When comparing privacy norms and zero-shot, we observe that providing privacy norms to the LLMs can improve appropriate judgments for the SPA dataset across all models, and improve inappropriate judgments for the Education dataset on most models. However, including general norms can sometimes degrade performance. For instance, with Gemini 2.5 Pro on the Education dataset, $F1_{\text{A}}$ decreases by $22\%$ while $F1_{\text{I}}$ improves by only $10\%$, suggesting that adding general privacy norms is not always beneficial.
(5) Prior user privacy judgments improve personalized privacy judgments over general privacy norms. By comparing the $F1_{\text{A}}$ and $F1_{\text{I}}$ between ICL and both zero-shot and privacy norms for all datasets and models, it is clear that personalizing LLMs with prior user privacy judgments improves appropriate and inappropriate judgments over general privacy norms. In some cases, the improvement over general privacy norms can be as large as 50% for $F1_{\text{A}}$.
(6) Personalized privacy judgments improve when including undetermined as an option. When comparing ICL and ICL (w Undet), we observe an improvement in $F1_{\text{A}}$ and $F1_{\text{I}}$ for all LLMs. These results suggest that giving an LLM the option to abstain reduces the likelihood that the model forces an erroneous appropriateness judgment.
6.2. Ablation Evaluation
We evaluate how the number of prior requests affects the performance of ARIEL and baseline approaches. Moreover, we investigate how the choice of reasoning affects LLM privacy judgments.
Figure 6. Ablation study on the number of prior requests with user judgments contained in user memory $D_{u}$. We evaluate ICL (w Undet) and ARIEL on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class, and the total number of appropriateness judgments (Sup). We find that ARIEL is more robust to varying the number of prior requests compared to ICL w (Undet).
Figure 7. Ablation study on CoT. We evaluate zero-shot, privacy norms, and ICL w (Undet) on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for the appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class. The results suggest that generally CoT has negligible impact on privacy judgments across both models.
6.2.1. Number of prior requests
We explore how varying the number of prior requests with user judgments affects the agent’s judgments. Specifically, we compare the $F1_{A}$, $F1_{I}$, and Support (Sup) of ARIEL against ICL (w Undet) on the SPA dataset using Gemini 2.5 Pro. As evident in Figure 6, while the $F1_{I}$ gap remains relatively constant, the disparity in $F1_{A}$ widens significantly as the number of prior requests decreases. Notably, when reducing the prior requests from 60 to 20, $F1_{A}$ drops by only 3.4% for ARIEL, compared to an 8.8% decline for ICL (w Undet). This demonstrates that ARIEL is more robust to limited prior privacy judgments. Nonetheless, we observe that the number of entailments generated by ARIEL decreases substantially as the number of prior requests decreases. This result is expected, as there are fewer candidate requests available to satisfy the entailment criteria (i.e., fewer prior requests differ by exactly one parameter from the incoming request).
6.2.2. Assessing the impact of reasoning with Chain of Thought (CoT)
To assess the impact of reasoning strategy choice on appropriateness judgments, we implement CoT by prompting the LLM to provide a reasoning before making a judgment (wei2022chain; yang2023large). Inspired by findings that CoT affects privacy leakage (mireshghallah2024can), we evaluate its effect on judgment accuracy across zero-shot, privacy norms, and ICL (w Undet). Figure 7 indicates that CoT typically provides negligible changes in performance. However, we observe for zero-shot with Gemini 2.5 Pro that the $F1_{A}(\%)$ and $F1_{I}(\%)$ are significantly impacted by CoT, both positively and negatively.
While CoT provides mixed results for general privacy norms, the results clearly demonstrate that adding CoT to both ICL and ICL (w Undet) consistently hurts their performance. When manually inspecting the generated reasoning traces, we notice that the LLM makes misaligned assertions about the user’s privacy judgments, which propagate all the way to the model’s final judgment (see Section 6.3 for more information). This result also demonstrates that eliciting explanations from LLM-based reasoning can degrade appropriateness judgments. In contrast, ARIEL provides a clear, traceable reasoning while maintaining well-grounded personalized privacy judgments.
6.3. Qualitative Analysis
6.3.1. Error Analysis on Appropriateness Judgments for ICL (w Undet) + CoT
Analysis of the reasoning traces reveals that Gemini 2.5 Pro occasionally performs implicit entailment when identifying prior requests that share identical parameters except for one. This behavior is noteworthy, as we did not explicitly instruct the model to use this logic, but only to extract patterns from user privacy judgments. We find that the majority of errors stem from preference misalignment (14 examples), where the model draws incorrect equivalencies between distinct parameter values to perform this entailment. For instance, given prior requests involving ”close friends” or ”parents” as recipients and an incoming request involving ”housemates,” the model infers that the privacy judgment should remain consistent due to perceived similarity in recipient type. However, the user may not view these three recipients as similar.
We also attribute a subset of errors to the model’s inherent conservatism regarding information sharing, which we describe as default to inappropriate (5 examples). This occurs when the model concludes it cannot use the most similar prior requests that are judged as ”appropriate” by the user. However, rather than outputting ”undetermined”, the model reasons that the judgment is inappropriate. Finally, we identified one instance of contradictory prior requests, where the provided user history contained judgments that logically conflicted with the ground truth for the incoming request.
@@ 6.3.1. Error Analysis on Appropriateness Judgments for ICL (w Undet) + CoT
6.3.2. Error Analysis on Appropriateness Judgments for ARIEL
@@
We classify errors in ARIEL into three categories: (1) ontology generation, (2) ontology mapping, and (3) contradictory prior requests. The majority of errors stem from ontology generation, where the generated ontology produces incorrect entailments (10 examples). This typically arises from data sparsity when specific parameter values (e.g., ’housemates’ or ’close friends’) appear only once or twice in the user’s prior requests. The model struggles to accurately reflect the user’s privacy preferences onto the ontology. A second source of error is ontology mapping, where the model fails to map parameter values to the ontology correctly (5 examples). This is often caused by domain-specific idiosyncrasies in the evaluation dataset. For instance, the transmission principle has a specific format combining both purpose and condition, which can be challenging for the model to map. Finally, the remaining errors (5 examples) result from contradictory prior requests. For instance, there was a prior request where the user had judged sharing bank account details with business and finance skills for the purpose of making transactions as inappropriate. The incoming request is identical to the prior, except there is no purpose. Thus, the incoming request should also be inappropriate, since removing the purpose likely makes the request more inappropriate. Both ICL (w Undet) + CoT and ARIEL judged this request inappropriate; however, the user judged the incoming request to be appropriate.
Potential solutions for (1) involve providing more user judgments on prior requests during ontology generation, while solutions for (2) involve providing more information on the domain-specific format of each parameter in the ontology mapping prompt. The errors in (3) represent an open challenge that is discussed in more detail in Section 7.1.
@@ 7. Discussion
Based on our results and takeaways, we discuss in more detail the broad implications of this work as well as its limitations.
7.1. Personalization Challenges
In this paper, we consider the user’s judgments on prior data-sharing requests as proxies for the user’s privacy preferences (wijesekera2017feasibility; colnago2022concern). Personalizing privacy decisions based on these preferences introduces several challenges that must be considered.
Contradictory Privacy Preferences
Highlighted by the error analysis in Section 6.3, contradictions can arise within the user’s own history of prior judgments. For example, the user says it is inappropriate to share their banking information with a third party if their data was confidential, but deems it appropriate if there is no safeguard on the data. In such scenarios, the agent may execute the entailment logic correctly but still reach an incorrect privacy judgment. Consequently, the agent must (1) detect these user privacy judgment contradictions and (2) resolve them, typically by escalating the request to the user. The fundamental difficulty lies in distinguishing whether an incorrect privacy judgment stems from an error in the entailment process of ARIEL or an inconsistency in the user’s prior privacy judgments.
Evolving Privacy Preferences
The second challenge involves the dynamic nature of privacy preferences as they can change over time. For example, a user can become more conservative about sharing their data as they get older. Currently, ARIEL does not support automatic detection of privacy preference changes. Hence, future adaptations of ARIEL require a mechanism to detect when the user’s privacy preferences are shifting and determine which preferences must be updated. One potential mechanism is to periodically escalate incoming requests to the user to confirm the agent’s privacy decision.
7.2. User Agency in Personal AI Agents
While personalized agents reduce the burden of privacy management, they must not remove user agency (mitchell2025fully; zhang2025autonomy). This observation motivated the entailment logic of ARIEL, where the agent automates clear decisions that can be entailed but escalates requests where the judgment cannot be easily inferred by the user’s prior judgments. Our results confirm that identifying these non-entailed incoming requests drives ARIEL’s judgment accuracy gains. Furthermore, by providing transparent reasoning traces for user auditing, ARIEL successfully balances automation with necessary human oversight.
Although ARIEL was designed with user agency as a core requirement, further work is needed to strengthen the interaction between the user and the agent. One promising direction is the development of a dedicated user interface (UI) that enables users to: (1) visualize the reasoning steps behind ARIEL’s decisions (e.g., which requests were entailed and how they were mapped in the ontology), and (2) intervene in the decision-making process when necessary. Subsequently, this UI can facilitate user studies to evaluate how well ARIEL’s privacy reasoning and judgments align with the user’s own privacy reasoning (nezhad2026understanding). Furthermore, this UI could be used to evaluate potential solutions for the challenges outlined in Section 7.1, such as user verification of ARIEL’s reasoning and decisions to identify contradictions or changes in the user’s privacy preferences.